The Blade Division of Dana Company produces hardened steel blades. Approximately one-third of the Blade Division's output is sold to the Lawn Products Division of Dana; the remainder is sold to outside customers. Blade Division's estimated sales and cost data for the year ending June 30th are as follows:

Sales to Lawn Products Division Sales to Outsiders
Revenue $ 28,500 $ 76,000
Variable costs 19,000 38,000
Fixed costs 4,800 19,500
Gross margin $ 4,700 $ 18,500
Unit sales 19,000 38,000

The Lawn Products Division has an opportunity to purchase, on a continual basis, 10,000 blades (of identical quality) from an outside supplier, at a cost of $1.70 per unit. Assume that the Blade Division cannot sell any additional products to outside customers. Assume, too, that there are no short-term avoidable fixed costs. Based solely on short-term financial considerations, should Dana allow its Lawn Products Division to purchase the blades from the outside supplier, and why?

Multiple Choice

A) Yes, because buying the blades would save Dana Company $5,000.

B) No, because making the blades would save Dana Company $6,000.

C) Yes, because buying the blades would save Dana Company $7,000.

D) No, because making the blades would save Dana Company $7,000.
Answer #1
Variable cost per unit of making 1 =19000/19000
Total cost of making 10,000 blades 10000 =10000*1
Total cost of buying 10,000 blades 17000 =10000*1.70
Additional cost of buying 7000 =17000-10000
No, because making the blades would save Dana Company $7,000.
Option D is correct